Identify the greater number, wherever possible, in each of the following?
28 or 82

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

The expansion of 28= 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 = 256
The expansion of 82 = 8 × 8= 64
Clearly,
256 > 64
So, 28> 82
Hence 28 is thegreater number.

Thanks